OpenStudy (anonymous):

Compute lim xtan (1/x) as x approaches infinity Compute lim xtan (1/x) as x approaches infinity @Mathematics

OpenStudy (anonymous):

its an inderterminate product right?

OpenStudy (zarkon):

let u=1/x take limit as\[u\to0^+\]

OpenStudy (zarkon):

use \[\lim_{u\to 0}\frac{sin(u)}{u}=1\]

OpenStudy (anonymous):

why is it sin (u) over u

OpenStudy (anonymous):

it's not it's sin(u)/u * cos(u) = 1 * 1

OpenStudy (anonymous):

\[\lim_{u \rightarrow 0}\frac{1}{u}\frac{sinu}{cosu}=\lim_{u \rightarrow 0}\frac{1}{cosu}*\lim_{u \rightarrow 0}\frac{sinu}{u}\]

OpenStudy (anonymous):

whoops, 1/cos(u)...yeah

OpenStudy (anonymous):

both limit equal one so 1*1=1

OpenStudy (zarkon):

just something to note \[\lim_{x\to\infty}x\tan(1/x)\] is equivalent to \[\lim_{u\to0^+}\frac{\tan(u)}{u}=\cdots\] but showing \[\lim_{u\to0}\frac{\tan(u)}{u}=1\] will technically work.

OpenStudy (anonymous):

so it wasnt an indetermiantae form since it equaled 1? or was it an indeterminate product?

OpenStudy (zarkon):

it is an indeterminate form